Circular No. 6579

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 55P/1997 E1 (TEMPEL-TUTTLE) O. R. Hainaut, University of Hawaii, reports his measurements of the

recovery images of the comet associated with the Leonid meteors, as follows:

1997 UT α_{2000} δ_{2000} m_{2} Observer
Mar. 4.62838 13^{h}43^{m}12.^{s}63 +4^{o}43′37′′.7 22.5 Meech
4.637 13 43 12.31 +4 43 41.7 _{”}
4.64876 13 43 11.57 +4 43 49.0  _{”}
4.65752 13 43 11.07 +4 43 54.4  _{”}
7.3108 13 40 40.80 +5 10 45.3 22  Martin

K. J. Meech, O. R. Hainaut and J. Bauer (Mauna Kea). Keck II 10-m reflector + LRIS CCD. Some interference from clouds, but comet clearly visible as a point source. No coma; image profile similar to stars. Seeing 0′′.8. Magnitude is crude R estimate using 5′′ diaphragm.

P. Martin, K. Mueller, G. Van de Steene, N. Hurtado and J. Miranda (European Southern Obs.). 3.6-m New Technology Telescope + SuSI. Comet clearly visible on superb 15-min exposure. No coma. Seeing 0′′.7.

Two recent predictions for comet 55P/Tempel-Tuttle have been made using the observations in 1865–1866, supplemented by the very scanty data from 1366, 1699 and 1965, the comet having been recovered on that last occasion following an insightful prediction by J. Schubart (IAUC 1907, 1926 and 1979). The prediction by S. Nakano and I. Hasegawa (MPC 27288), giving T = 1998 Feb. 28.079 (at epoch 1998 Mar. 8), requires correction by ΔT ∼ −0.14 day; that by D. K. Yeomans et al. (1996, Icarus 124, 407), giving T = 1998 Feb. 28.091, requires correction by ΔT ∼ −0.06 day. For the computation of the following orbital elements, fitted to the nine observations in 1965 and 1997, the Yeomans value of e was assumed:

Epoch = 1998 Mar. 8.0 TT

T = 1998 Feb. 27.977 TT  ω = 172.^{o}484
e = 0.90551  Ω = 235.251 }2000.0
q = 0.97657 AU  i = 162.485

a = 10.33467 AU n^{o} = 0.029666 P = 33.22 years

1997 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{2}
Mar. 3 13^{h}44.^{m}67 + 4^{o}27.^{′}6 3.678 4.463 137.^{o}8 8.^{o}6 22.7
13 13 34.78 + 6 10.7 3.491 4.375 149.4 6.7 22.4
23 13 22.77 + 8 02.3 3.338 4.286 159.5 4.7 22.2
Apr. 2 13 08.99 + 9 56.7 3.227 4.197 164.1 3.8 22.0
1997 March 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden